OpenStudy (anonymous):

1. Solve 3x − 9 + x > 27 a. x > 4.5 b. x < 4.5 c. x > 9 d. x < 9 3. Maria earns $8 an hour babysitting. Last week Maria made $64. She needs to have at least $150 for vacation. Which inequality can be used to determine how many more hours Maria needs to babysit before she is ready for vacation? a.64x + 8 ≥ 150 b. 8x + 64 ≥ 150 c. 64x + 8 ≤ 150 d. 8x + 64 ≤ 150 5. Solve 4(5m + 6) − 8 ≤ 336. a. m ≤ 16.9 b. m ≤ 16 c. m ≥ 16.9 d. m ≥ 16 6. Solve −4x + 15 > 59. a. x < 11 b. x > −11 c. x < 11 d. x < -11

OpenStudy (anonymous):

I lined up number 5 wronggggg,

OpenStudy (anonymous):

\[3x − 9 + x > 27 \] combine like terms \[4x-9>27\] add \(9\) \[4x>36\] divide by \(4\) \[x>9\]

OpenStudy (anonymous):

On that one, don't you change the sign because you divide?

OpenStudy (anonymous):

Like, it wouldn't be x >9 it would be x < 9 because you divided, yeah?

OpenStudy (anonymous):

\[4(5m + 6) − 8 ≤ 336.\] add 8\[4(5m+6)\leq 344\] divide by \(4\)\[5m+6\leq86\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no

OpenStudy (anonymous):

only change the inequality if you multiply or divide by a NEGATIVE number 4 is positive

OpenStudy (anonymous):

Right, right! Thanks! Could you help me on the rest?

OpenStudy (anonymous):

this one \[−4x + 15 > 59. \] you would change the sign

OpenStudy (anonymous):

\[-4x+15>59\] subtract 15\[-4x>44\] divide by \(-4\) and now you have to change it becuase \(-4\) is negative \[x<-11\]

OpenStudy (anonymous):

you can solve this \[5m+6\leq 86\] right?

OpenStudy (anonymous):

I didn't understand number 5? /:

OpenStudy (anonymous):

i wrote the solution above, did you see it? or would you like me to repeat it?

OpenStudy (anonymous):

lol, don't be rude, I would divide 86 to 4? I don't understand

OpenStudy (anonymous):

\[ 4(5m + 6) − 8 ≤ 336.\] add 8\[4(5m+6)\leq 344\]divide by \(4\) \[5m+6\leq 86\] subtract \(6\)\[5m\leq 80\] divide by 5\[x\leq 16\]

OpenStudy (anonymous):

OHHHHH! Damn, I see!

OpenStudy (anonymous):

kk that is all right?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

did 3 first

OpenStudy (anonymous):

ooh i see, did 1 first 3 is not done yet

OpenStudy (anonymous):

at least means "greater than or equal to" as in "you have to be at least 21 to drink in this bar"

OpenStudy (anonymous):

she already has 64 dollars, and makes $8 per hour, go with \[8x+64\geq 150\]

OpenStudy (anonymous):

Awesome! :)

OpenStudy (anonymous):

awesome is not math on saturday night

OpenStudy (anonymous):

Omg, I know :(

OpenStudy (anonymous):

How would you put this iquality? If you add 13 to four times a certain number, the sum is less than 265. Let x represent the number. Write an inequality that can be solved to find the value of x.

OpenStudy (anonymous):

@satellite73